• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name jpchen

    Log entry time 03:01:41 on December13,2007

    Entry number 220310

    keyword=Raster size changed (by itself!) during 12/11 swing

    Raster size change happened on Monday (12/10) swing shift sometime between run 3684 (21:46-21:50) and run 3686 (22:11-22:15). Clearly raster size changed, probably due to beam tune change. However, there is no reported change of beam tuning at that time. During that time, there was a Loop3 excursion and target vacuum rising
    (Alog 219736/219737). Elog entry show the Hall C current increase (Elog #1406643) (also some interlock change removal). Not sure if
    any of the above is related.


    FIGURE 1

    FIGURE 2